OpenStudy (anonymous):

find the distance between two cities if train running between two cities arrives at its destination 10 minutes late when it goes 40 miles per hour and 16 minutes late when 30 miles per hour

OpenStudy (anonymous):

how to find Integration of p2+p-2 w.r.t. dx

OpenStudy (dan815):

v = d/t so distance = v*t the first equation is d=40*(t+10) can u think about the 2nd equation

OpenStudy (raden):

first, solve for t 40(t + 1/6) = 30(t + 4/15)

OpenStudy (raden):

after u get the value of t, then subtitute t to one of equations 40(t + 1/6) or 30(t + 4/15)